OpenStudy (joftheworld):

logarithmic problems .. @satellite73 help please ..

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

3 ducks and 5 pelicans

OpenStudy (anonymous):

\[\log(A)+\log(B)=\log(AB)\] so \[\log(11)+\log(x)=\log(11x)\]

OpenStudy (jdoe0001):

http://www.chilimath.com/algebra/advanced/log/images/rules%20of%20exponents.gif <--- use the 1st rule listed there

OpenStudy (joftheworld):

thats it? just 11x ?

OpenStudy (joftheworld):

i thought there would be more to it..

OpenStudy (joftheworld):

what about this..

OpenStudy (anonymous):

lol stick a 17 in the box what else could it be?

OpenStudy (joftheworld):

i figured lol how about this ?

OpenStudy (anonymous):

\(2\times 9=18\) goes out front

OpenStudy (joftheworld):

thanks !